OpenStudy (goggles):

find the sum of the first 20 terms of the arithmetic sequence 4, 10, 16, 22....

OpenStudy (anonymous):

Well, on thing you should notice is that each number is changing by a value. What is that value?

OpenStudy (anonymous):

one*

OpenStudy (goggles):

6

OpenStudy (anonymous):

Good. Now, we can find the 20th term in the sequence by using the formula\[a _{n}=a _{1}+(n-1)d\] We know n=20, and we know a_1=4 and you just said d=6, so just plug all those numbers in and you can find the 20th term. Then you can find the sum of the first 20 terms by using the formula \[S _{n}=\frac{ n }{ 2 }\left[ a _{1} +a _{n}\right]\]

OpenStudy (goggles):

thanks

OpenStudy (anonymous):

You are welcome!